OpenStudy (anonymous):

I have an integral from t=0 to 40 of (t/40)^(a+1)dt. I can get have way to the known answer of 40/(a+2) by substitution to using u=(t/40) and du=(1/40) which resolve to 40 times the integral over 0 to 1 of u^(a+1)du. I am simply stumped now how to get to 40/(a+2). What rules am I not seeing? Cheers A.

hartnn (hartnn):

you know the formula ., \(\Large \int x^n dx = \dfrac{x^{n+1}}{n+1}+c\)

hartnn (hartnn):

using that, \(\Large 40\int u^{a+1}du=40 \dfrac{u^{a+2}}{a+2}\)

hartnn (hartnn):

then apply your limits of u=0 and u=1

OpenStudy (anonymous):

Okay, thanks. I will have a scratch at that and see how it gels. I am doing my masters and unfortunately haven't used calculus in 20+ years. Fun but slow.

OpenStudy (anonymous):

DOH! One raised to any power is one. So close but sooooo rusty. Thanks hartnn.

hartnn (hartnn):

welcome ^_^ oh, and since you are new here, \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{E}}\color{#e6e600}{\text{L}}\color{green}{\text{C}}\color{blue}{\text{O}}\color{purple}{\text{M}}\color{purple}{\text{E}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{n}}\color{orange}{\text{S}}\color{#e6e600}{\text{t}}\color{green}{\text{u}}\color{blue}{\text{d}}\color{purple}{\text{y}}\color{purple}{\text{!}}\color{red}{\text{!}}\color{orange}{\text{ }}\color{#e6e600}{\text{:}}\color{green}{\text{)}}\color{blue}{\text{}}\end{array} \]